OpenStudy (anonymous):

The Garners rented a moving van for $59.95 plus $0.50/mi. Before returning the van, they filled the gas tank, which cost $21.50. The total cost for renting the van, including gas was $199.45. To the nearest mile, how many miles did the Garners drive the van?

OpenStudy (anonymous):

First subtract the price of gas:\[199.45-21.5=177.95\]Then the base price of renting the moving van:\[177.95-59.95=118\]Then divide by the price per mile:\[118/.5=118/(1/2)=118*(2/1)=118*2=236\] The Garners drove the van 236 miles.